### v2

# HAIFIT: Human-Centered AI for Fashion Image Translation

Abstract
In the realm of fashion design, sketches serve as the canvas for expressing an artist’s distinctive drawing style and creative vision, capturing intricate details like stroke variations and texture nuances. The advent of sketch-to-image cross-modal translation technology has notably aided designers. However, existing methods often compromise these sketch details during image generation, resulting in images that deviate from the designer’s intended concept. This limitation hampers the ability to offer designers a precise preview of the final output. To overcome this challenge, we introduce HAIFIT, a novel approach that transforms sketches into high-fidelity, lifelike clothing images by integrating multi-scale features and capturing extensive feature map dependencies from diverse perspectives. Through extensive qualitative and quantitative evaluations conducted on our self-collected dataset, our method demonstrates superior performance compared to existing methods in generating photorealistic clothing images. Our method excels in preserving the distinctive style and intricate details essential for fashion design applications.  [^1] Code is available at https://github.com/ExponentiAI/HAIFIT/.
Index Terms:
Sketch-to-Image Generation; Image-to-Image Translation; Generative Adversarial Networks
I Introduction
In the creative journey of design, sketches serve as the foundational tool for conceptualizing and outlining new pieces, demonstrating the unique styles of different designers. Designers initiate their process by crafting sketches that capture the basic structure of their envisioned creation. Subsequently, through iterative refinement, these sketches evolve, receiving modifications and intricate detailing to further refine the designs. Progressing from these refined stages, designers infuse life into their sketches by adding colors, aiming to impart depth and realism to their representations. Throughout this process, a robust sketch-to-image method plays a crucial role. It not only allows designers to preview renderings of colored sketches, aiding in refining the sketchwork, but also facilitates the generation of diverse image styles to inspire their designs.
In the earlier research, methods like Sketch2Photo [4] synthesized realistic images by amalgamating object and background images derived from provided sketches. They laid the groundwork for deep learning techniques. The emergence of Generative Adversarial Networks (GAN) [6] marked a significant shift in this domain, with methods such as BicycleGAN [35], pix2pix [12], and others [11, 18, 15, 31] gaining prominence. Although these methods have showcased notable results, preserving sketch details remains a challenge, as illustrated in Fig. 1. Through our comparison of two classic algorithms based on GAN [29] and Diffusion Model (DM) [17], we discovered significant hurdles in retaining sketch details. Specific elements such as clothing necklines, pockets, sleeves, and other intricate details prove challenging to effectively express in the lines of the designer’s sketches. This limitation significantly hampers designers’ ability to preview and enhance finer details and refinements within sketches.
Figure 1: The performance of different methodological approaches [29, 17] compared to ours in handling sketch details (zooming in for better visual).
In recent years, with the introduction of the diffusion model [9], a novel and improved image generation algorithm has swiftly emerged. Some works, such as DDIM [26], employ skip sampling to expedite the model’s inference process, while LDM [22] utilizes feature mapping to execute the diffusion process in the latent space, effectively mitigating the issue of high resource requirements for model training. Additionally, Palette [24] and BBDM [17] showcase the efficacy of diffusion models in image translation tasks. However, these methods heavily rely on extensive data for learning and demand substantial computing resources. Consequently, they still face significant challenges in model training, inference, and fine-tuning, resulting in high time and resource costs, which may not be user-friendly for designers.
To address these challenges, we introduce HAIFIT (Human-Centered AI for Fashion Image Translation), a GAN-based creative generation network designed for sketch-to-image translation in fashion design. HAIFIT aims to empower designers to foster inspiration by preserving as much detail as possible from their sketches in the generated images. Specifically, HAIFIT focuses on converting clothing sketches into high-quality realistic clothing images using a pyramidal generative adversarial network. To facilitate this conversion, we have developed a customized Multi-scale Feature Fusion Encoder (MFFE) that maps the reference hand-drawn sketches and generated images into a shared latent space. MFFE comprises a Shallow Convolution Module (SCM) and an Abstract Feature Representation Module (AFRM), which enables the capture of long-term dependencies between different regions and improves the global correlation in feature representations. Furthermore, we enhance the fine-grained features among generated images of various scales by integrating the Cross-level Skip Connection Module (CSCM) within the feature space across different layers. By employing multi-scale discriminators, we ensure the preservation of global structural integrity while emphasizing fine details, resulting in more realistic generated images. The key contributions of our work are summarised as follows:
  • Methodology. To the best of our knowledge, we are the pioneers in treating sketch features as sequence inputs, ensuring consistency by inputting features from both forward and reverse feature sequences. This innovative approach enhances the representation of details in the generated images.
  • Structure. We introduce the LSTM-based MFFE to enhance the representation of global features and ensure coherence and consistency in the latent space. Additionally, our CSCM strategically directs the model’s attention to finer sketch details, contributing to improved and more nuanced sketch feature representation.
  • Datasets. We construct a background-free dataset comprising fashion clothes and their corresponding sketches for cross-modal image translation tasks, filling the gap in this field.
  • Performance. Our method demonstrates superior performance in extensive quantitative and qualitative experiments compared to various baselines.
II Related Work
II-A Sketch-to-Image Generation
The task of sketch-to-image generation falls within the realm of image translation, a pivotal objective in computer graphics and computer-human interaction. In the domain of fashion clothing design, having an effective algorithm that can accurately generate diverse clothing images based on designers’ sketches is invaluable. Such algorithms enable designers to express their creative ideas effectively, save development time, and enhance design efficiency. Early methods, like Photosketcher [5], utilized local features for image retrieval and employed graphics techniques to segment and stitch images together to achieve sketch-to-image generation. Subsequently, with the development of convolutional neural networks [16, 27, 25, 7], deep learning-based methods have become the mainstream approach for achieving sketch-to-image generation. Isola et al. [12] introduced a unified framework for image-to-image translation models adaptable to various scenarios. This framework facilitates mapping segmented images to natural images or converting edge maps to color images, among other applications. Building upon this, Pix2piHD [29] utilized multi-scale generators and discriminators to achieve high-resolution image translation. Several other methods [11, 1, 18, 15] have improved image generation performance by incorporating attention networks, innovative adversarial loss functions, or different image features. However, these methods still encounter challenges in handling complex strokes and refining synthetic structures to produce realistic images. In summary, synthesizing an image that faithfully represents the content of a sketch while maintaining consistent styles is paramount.
II-B Generative Adversarial Networks
The Generative Adversarial Network (GAN) comprises two components: the generator and the discriminator. It leverages adversarial training to enhance the generator’s ability to generate realistic images, yielding significant achievements in image generation, restoration, and various other domains. Initially proposed by Goodfellow et al. [6]. For generating realistic data, GANs have since undergone several extensions and improvements. Mirza et al. [20] extended the GAN to enable the generation of images with specific conditions, allowing users to control attributes such as image categories. Radford et al. [21] introduced convolutional neural networks into both the generator and discriminator, thereby enhancing GAN’s ability to capture spatial image features. Zhu et al. [34] achieved cross-domain image conversion by introducing a cycle consistency loss function, enabling realistic transformations from one domain to another, such as from horse to zebra. Karras et al. [14] achieved stable control over the style and diversity of generated images by introducing style variables. Brock et al. [3] further scaled up the GAN and improved training methods to generate high-resolution images. Additionally, methods like WGAN [2], PGGAN [13], and SAGAN [32] have introduced new optimization losses, multi-layer resolution training strategies, and self-attention mechanisms to achieve more stable GAN training and further enhance the quality and diversity of generated images. In general, the continuous development and evolution of GANs have unlocked significant potential in the field of image generation, transforming the generation of realistic images from a concept into reality. Compared to the recent diffusion models [24, 17], GANs possess the advantages of low hardware costs and fast synthesis speed, thus continuing to play a key role in the design field.
III Methodology
Algorithm 1 The PGGAN Training Process.
0:  $n$: Number of multi-layers
0:  $G=\{G_{1},G_{2},...,G_{n}\}$: Generator networks
0:  $D=\{D_{1},D_{2},...,D_{n}\}$: Discriminator networks
0:  $I=\{I_{1},I_{2},...,I_{n}\}$: Image resolutions
0:  $k$: Number of training iterations
1:  $G\leftarrow G_{1}$, $D\leftarrow D_{1}$, $I\leftarrow I_{1}$
2:  Initialize $G_{1}$ and $D_{1}$ with random weights
3:  for $i=1$ to $n$ do
4:     Training $D_{i}$ for $k$ iterations with fixed $G_{i}$
5:     Training $G_{i}$ for $k$ iterations with fixed $D_{i}$
6:     if $I_{i}<I_{n}$ then
[…]
8:        Initialize new layers $G_{i+1}$ and $D_{i+1}$ for higher resolution
9:        Update weights of existing layers in $G$ and $D$
10:     end if
11:  end for
In this study, we present HAIFIT, as depicted in Fig. 2, which is based on PGGAN [13] and comprises two core components: the Multi-scale Feature Fusion Encoder (MFFE), and the Cross-level Skip Connection Module (CSCM). The MFFE is used to get global contour features and abstract intent features. Then the feature maps are resized and concatenated at multiple scales before being fed into the multi-scale pyramid generator, which can produce realistic images with complete structures and exquisite details.
III-A Preliminaries
The traditional GANs [6] enhance the model’s generation capabilities by adversarially training the generator and discriminator. However, they still face significant challenges in terms of image generation quality and training process stability. To address these issues, PGGAN [13] introduces a progressive growth training method. This approach starts from low-resolution images and gradually increases the number of layers and image resolution of both the generator and discriminator. This progressive growth training process not only simplifies the model’s convergence to generate higher-quality and more flexible-resolution images, but also mitigates issues such as mode collapse and gradient disappearance during the training process. Consequently, we utilize the PGGAN as the framework for our approach.
The specific algorithmic process is outlined in Algorithm 1. We commence training the initial generator $G_{1}$ and initial discriminator $D_{1}$ from the initial resolution $I_{1}$, gradually increasing the resolution and continuing training by adding new $G_{i+1}$ and $D_{i+1}$ layers. This process continues until the final resolution $G_{n}$ and $D_{n}$ model training is completed, ultimately yielding the desired generated image.
Figure 2: The illustration of our HAIFIT architecture. Given $M$ paired sketch-image samples $\{(x_{i},y_{i})\in(X,Y)\}^{M}_{i=1}$, our goal is to map images from hand-drawn sketches domain $X$ to realistic images domain $Y$, while preserving as many line details from the sketch $X$ as possible in the generated image $\overline{Y}$.
III-B Multi-scale Feature Fusion Encoder
The Multi-scale Feature Fusion Encoder (MFFE) consists of two modules: the Shallow Convolutional Module (SCM) and the Abstract Feature Representation Module (AFRM). As illustrated in Fig. 3, the SCM employs a convolutional network to learn essential global contour information, such as the structure of the object. Simultaneously, inspired by the LSTM architecture  [10], we design the AFRM to learn long-range dependencies in key areas of the input sketch from various perspectives, which can effectively capture style features and geometric features of strokes. By combining these features, we can effectively capture the consistency of objects’ complete structure and intricate textures within the latent space.
As shown in Fig. 2, in the bottom layer, only the sketch features are inputted. In the other layers, we perform element-wise addition between the generated image features and the sketch features, and then input the resulting summed features into the MFFE. In the SCM we employ a sequence of convolutional layers, denoted as $Con$, to extract the global contour features and shape information $\widetilde{x}$ from the input sketch $x_{s}$. Here, $x_{s}\in\mathcal{R}^{b\times 3\times m\times m}$, $\widetilde{x}\in\mathcal{R}^{b\times 256\times m/4\times m/4}$, $b$ is the batch size, and $m$ depends on the network layer, so we can express this process as $\widetilde{x}=Con(x_{s})$. Specifically, the SCM extracts features in a blockwise manner, which will disrupt the continuity of strokes and result in pixel-level contour features. Since sketches are continuous and composed of multiple lines in different directions, it is crucial to consider the correlation between adjacent lines in order to effectively learn the line features of complex sketches.
To address this issue, we introduce the AFRM to learn the feature $\widetilde{x}$. The AFRM consists of a two-layer LSTM network, capturing abstract drawing intent features from four directions. After downsampling $\widetilde{x}$, we obtain the feature $\widetilde{x^{\prime}}\in\mathcal{R}^{b\times 256\times 4\times 4}$. We then swap feature positions after merging its second and third dimension features to obtain $\mathcal{R}^{4\times b\times 1\times 1024}$ and divide it into four parts, namely $\widetilde{x}_{0}$, $\widetilde{x}_{1}$, $\widetilde{x}_{2}$, and $\widetilde{x}_{3}$, with dimensions $\mathcal{R}^{b\times 1\times 1024}$. These features are then individually sequence inputted into the LSTM network, generating features $\overline{x}_{0}$, $\overline{x}_{1}$, $\overline{x}_{2}$, and $\overline{x}_{3}$ by recursively processing the input features from different directions. Next, we concatenate the features $\overline{x}_{0}$, $\overline{x}_{1}$, $\overline{x}_{2}$, and $\overline{x}_{3}$ to obtain the feature $\overline{x^{\prime}}\in\mathcal{R}^{4\times b\times 1\times 1024}$. Afterward, we restore the dimension of $\overline{x^{\prime}}$ to the dimension of $\widetilde{x^{\prime}}$. Finally, we apply a 1$\times$1 convolutional layer along the channel dimension to obtain the output feature $\overline{x}\in\mathcal{R}^{b\times 256\times 4\times 4}$.
The MFFE is employed to learn features from different perspectives of the sketches, which enables us to capture variations in lines and abstract painting intentions in crucial areas of the sketches. In the task of sketch-to-image generation, it is vital for the encoder to acquire diverse line patterns and painterly intent in clothing patterns and details. Failure to do so can severely impact the visual quality of the synthesized image, even if the overall generation quality is satisfactory. As a result, to further ensure the consistency of the generated images, we introduce the concept of reverse sketch feature input. This is because a sketch should yield consistent results regardless of the direction it is interpreted from. Therefore, the aforementioned process can be summarized as follows:
$\begin{split}\overline{x}_{dir}^{\prime}&=Cat(LSTM[\widetilde{x}_{0},\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3}])\\ \overline{x}_{rev}^{\prime}&=Cat(LSTM[\widetilde{x}_{0},\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3}])\\ \overline{x}&=Conv_{1\times 1}(Cat(\overline{x}_{dir}^{\prime},\overline{x}_{rev}^{\prime}))\end{split}$ (1)
Finally, we fuse the global structure feature $\widetilde{x}$ and abstract intention feature $\overline{x}$ through a differentiable parameter $\zeta$ with the initial value of 1.0 to get the embedded sketch feature $x$:
$x=\zeta\widetilde{x}+\overline{x}$ (2)
Figure 3: An overview of the Multi-scale Feature Fusion Encoder (MFFE).
III-C Cross-level Skip Connection Module
In order to effectively capture intricate details in the key areas of the sketch, we employ a pyramid generative learning approach [13] with a Cross-level Skip Connection Module (CSCM). Our model comprises a pyramid of generators $G_{n}\in\{G_{1},G_{2},\dots,G_{N}\}$, which are trained using a set of multi-scale images $x_{i}$ represented as ${x^{1}_{i},x^{2}_{i},\dots,x^{N}_{i}}$. Each generator $G_{n}$ is responsible for generating a realistic image $\overline{y}^{n}_{i}$. By adversarial training, the generator $G_{n}$ aims to deceive the associated discriminator $D_{n}\in\{D_{1},D_{2},\dots,D_{N}\}$, and the $D_{n}$ is trained to distinguish between the generated realistic images and real images. The generation of a realistic image sample starts at the coarsest scale $64\times 64$ and sequentially passes through all generators, progressing toward the finest scale $256\times 256$. At the coarsest scale, the MFFE takes a minibatch of the sketch feature maps $X_{n}$ as input. The effective receptive field at this level is typically half the height of the finer image. Each of the generators $G_{n}$ at finer scales adds details that are not generated by the previous scales. Therefore, in addition to the feature maps $X_{n}$, each multi-scale feature fusion encoder accepts an upsampled version of the generated image features from the coarser scale. Here, $X_{n}\in\mathcal{R}^{b\times 3\times m\times m}$, $Y_{n-1}\in\mathcal{R}^{b\times 3\times m/2\times m/2}$. This pyramid structure allows us to capture features at multiple scales and better handle the complex details of the sketch.
$Y_{n}=\begin{cases}G_{n}(X_{n}),&\mbox{if {n} = 1}\\ G_{n}(X_{n},Y_{n-1}),&\mbox{if {n} > 1}\end{cases}$ (3)
The generator $G_{n}$ consists of a residual block and multiple convolutional layers. In the coarsest layer, the residual block comprises four convolutional layers. In addition, in the pyramid network, as the number of pixels increases from bottom to top, the residual block progressively adds one convolutional layer at each level. This progressive increment facilitates the learning of sketch feature representations. Through numerous experiments, it has been observed that incorporating the CSCM between different layers in the pyramid network can enhance the quality of the generated images, which is illustrated in Fig. 4. In detail, We extract the output feature ${X^{\prime}}_{n-1}$ from the last layer of the residual block in generator $G_{n-1}$ and concatenate it with the output feature ${X}_{n}$ from the $n$-th layer of the MFFE. Subsequently, we use $\overline{X}_{n}$ as the input for generator $G_{n}$, where $\overline{X}_{n}\in\mathcal{R}^{b\times 256\times m\times m}$. In principle, the multi-level features of the same synthetic image should exhibit similarity. By employing cross-level feature operations, we can extract similar details from the features generated by different level generators and incorporate them as attention maps into the new feature map $X_{n}$. This facilitates image translation, resulting in more detailed and high-quality images.
$\begin{split}\overline{X}_{n}&=X_{n}+X_{n}\cdot X^{\prime}_{n-1}\\ Y_{n}&=G_{n}(\overline{X}_{n},Y_{n-1}),n=2,3,4\end{split}$ (4)
Figure 4: An overview of the Cross-level Skip Connection Module (CSCM).
III-D Loss Function
Our full loss function consists of L1 loss, adversarial loss, style loss, and perceptual loss. We empirically set the following parameters: $\lambda_{l1}$=1.5, $\lambda_{adv}$=10.0, $\lambda_{per}$=0.1, $\lambda_{style}$=250.0, to keep stable training and higher-quality results.
$\mathcal{L_{S}}=\lambda_{l1}\mathcal{L}_{l1}+\lambda_{adv}\mathcal{L}_{adv}+\lambda_{style}\mathcal{L}_{style}+\lambda_{per}\mathcal{L}_{per}$ (5)
L1 Loss. We directly employ the $L1$-distance loss to enhance the similarity between the pixel features of the generated image $G(x)$ and the real image $y$.
$\mathcal{L}_{l1}=\|y-G(x)\|_{1}$ (6)
Adversarial Loss. To improve the quality of generated samples, we utilize the vanilla GAN [6] min-max loss for the generator $G$ and discriminator $D$ training process.
$\mathcal{L}_{adv}=\mathbb{E}_{y}[\log D_{Y}(y)]+\mathbb{E}_{x}[\log(1-D_{Y}(G(x))]$ (7)
Style Loss. We incorporate the $Gram$ matrices-based loss to better reproduce the texture details of input sketches $x$.
$\begin{split}\mathcal{L}_{style}&=\|Gram(y)-Gram(G(x))\|_{1}\end{split}$ (8)
Perceptual Loss. To ensure both semantic similarity between the generated image and its ground truth, as well as encourage the network to generate diverse images, we employ the pre-trained VGG-16 [25] $\phi$ to extract $L$-levels features.
$\mathcal{L}_{per}=\sum_{i=1}^{L=5}\frac{1}{N_{i}}\|\phi_{i}(y)-\phi_{i}(G(x))\|_{1}$ (9)
IV Experiments
IV-A HAIFashion Dataset
Figure 5: Existing fashion datasets FashionGEN [23] and DeepFashion [19] compared to our newly proposed dataset, HAIFashion.
Considering our objective of learning clothing sketch features and generating high-quality clothing images while preserving the designer’s style and intricate details, it’s imperative to construct a new clothing dataset comprising paired images and sketches for model training and testing. While directly hand-drawing a large dataset of garment image-sketch pairs by designers may seem challenging, a more feasible approach involves designers creating a small amount of clothing image-sketch paired data, which can then be expanded using an image-to-sketch generation algorithm. Thanks to the personalized sketch generation algorithm proposed by StyleMe [31], this dataset construction becomes feasible. Thus, our dataset construction task focuses on building a fashion clothing image dataset for image-to-sketch generation.
Common publicly available fashion datasets like FashionGEN [23] and DeepFashion [19] predominantly feature realistic fashion images sourced directly from real-world photos. These images often include various backgrounds and individuals wearing the clothing in social settings. However, images with backgrounds pose significant challenges in generating single clothing sketches, which may not meet our requirements. The comparison between these datasets is illustrated in Fig. 5. In this study, we introduce a self-collected clothing dataset named HAIFashion. HAIFashion comprises 3100 pairs of fashion clothing image-sketch datasets. We achieved this by crawling 3100 background-free fashion clothing images from the internet and employing the algorithm [31] to generate designer-style clothing sketches, facilitating our experiments.
IV-B Implementation Details
We implement our model on Pytorch and all experiments are performed on a single NVIDIA GeForce RTX4090 GPU. We utilize the Adam optimizer with a learning rate of 1e-4 for the generator and 5e-4 for the discriminator, the batch size is set to 8. The learning rates decayed at the halfway point every 100 epochs. The training process continued until the SSIM [30] showed no improvement for 10 iterations. For the dataset, we use our proposed HAIFashion (§ IV-A) and divide it into 2500 (600) for model training and testing respectively.
TABLE I: The Overall sketch-to-image generation performance on the HAIFashion dataset. ($\uparrow$ indicates the higher value, the better effect. $\downarrow$ indicates the lower value, the better effect.)
Methods Venue Architecture Params Structural Perceptual
(MB) PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
BicycleGAN [35] NeurIPS’17 GAN 219.2 12.3482 0.4258 0.1976 70.6381
pix2pix [12] CVPR’17 GAN 45.6 19.2761 0.3826 0.1180 48.3830
MUNIT [11] ECCV’18 GAN 120.3 13.3265 0.4843 0.1947 74.2733
UAGITI [1] NeurIPS’18 GAN 2.4 11.8259 0.0901 0.2383 83.8309
pix2pixHD [29] CVPR’18 GAN 729.8 16.1863 0.5622 0.1026 48.0041
USTP [18] ECCV’20 GAN 31.4 13.2456 0.5126 0.1429 58.8756
UGATIT [15] ICLR’20 GAN 2764.8 12.8841 0.4799 0.1783 83.9190
Palette [24] SIGGRAPH’22 DM 250.7 7.8216 0.0787 0.3591 116.9303
BBDM [17] CVPR’23 DM 1,945.6 14.9788 0.5343 0.1676 59.7919
HAIFIT (Ours) - GAN 769.6 19.0660 0.6338 0.0778 28.5032
IV-C Evaluation Metrics
We employ four commonly used quantitative evaluation metrics:
  • PSNR (Peak Signal-to-Noise Ratio) assesses image quality by comparing the peak signal-to-noise ratio between the original image and the generated image. Higher PSNR values indicate smaller differences between the images and higher-quality generated images.
  • SSIM (Structural Similarity Index) [30] evaluates image quality by comparing the structural similarity between the original image and the generated image. It considers similarities in brightness, contrast, and structure to calculate the final score. Larger SSIM values signify higher similarity between the images and better generation effects.
  • LPIPS (Learned Perceptual Image Patch Similarity) [33] assesses image quality by utilizing a deep learning model and the L2 norm to calculate the perceptual difference between the generated image and the original image in the feature space. Lower LPIPS values suggest better-quality images.
  • FID (Fréchet Inception Distance) [8] evaluates image quality by comparing the statistical distribution in the feature space of the generated image with that of the original image. It employs the Fréchet distance to measure the similarity between the distributions. Lower FID values indicate higher-quality resulting images.
The first two metrics primarily gauge the completeness and structural similarity of the generated images, whereas the latter two metrics focus more on assessing the diversity, authenticity, and perceptual differences of the generated images.
IV-D Baselines
In our experiments, we compared our method with nine state-of-the-art baseline approaches. All models undergo identical preprocessing steps and are trained directly following the official open-source code to ensure a fair comparison.
  • BicycleGAN [35] focuses on modeling a distribution of possible outputs in a conditional generative model, enhancing the diversity of generated images.
  • pix2pix [12] implements image translation using conditional GAN to convert input images between different domains.
  • MUNIT [11] achieves image synthesis by decomposing image representations into content and style properties.
  • UAGITI [1] incorporates attention networks to minimize divergence between relevant parts of the data-generating distribution.
  • pix2pixHD [29] utilizes a novel adversarial loss and new multi-scale generator and discriminator architectures in conditional GAN for high-resolution image synthesis.
  • USTP [18] learns from unpaired sketch and photo data and tackles synthesis through shape translation and content enrichment stages.
  • UGATIT [15] employs an attention module and a learnable normalization function to enhance the quality and diversity.
  • Palette [24] proposes a unified framework for image-to-image translation based on conditional diffusion models, demonstrating the importance of self-attention.
  • BBDM [17] presents a novel image-to-image translation approach utilizing the Brownian bridge diffusion model, which enables direct learning of domain translation through a bidirectional diffusion process.
IV-E Comparison Results
Figure 6: The comparison of sketch-to-image generation from the baselines and our model.
Quantitative Experiments. We conducted quantitative evaluations to compare both the realism and faithfulness of the generated images. The quantitative results are summarized in Table I, despite achieving the second position in the PSNR index, it is important to note that this result is primarily due to the nature of PSNR, which emphasizes the signal-to-noise ratio and is sensitive to local noise and distortion in the image. HAIFIT, on the other hand, prioritizes retaining more details from the input sketches, including the initial outline, which may lead to a slight decrease in the PSNR index. However, overall, HAIFIT demonstrates excellent performance in all four metrics. This demonstrates the effectiveness of our method in sketch-to-image generation. This success can be attributed to the integration of our proposed MFFE and pyramid generator with CSCM. The MFFE helps our model learn object sketch contours and consider correlations between adjacent lines from different directions, enhancing stroke variation perception. Meanwhile, the pyramid generator with CSCM can capture specific details in key regions of complex sketches, facilitating the learning of the complete structure and fine-tuning of intricate details. The results confirm that our model can well adapt to sketches with specific detailed styles and generate more realistic images.
Qualitative Experiments. As shown in Fig. 6, it is evident that our proposed method outperforms other baselines by delivering high-definition and diverse images. For example, although BcycleGAN, USTP, and UGATIT produce some satisfactory results, the majority of their images exhibit a single color tone and lack fine details. Similarly, pix2pix, MUNIT, and pix2pixHD demonstrate good color diversity, but suffer from issues such as uneven color distribution and blurred details. UAGITI also faces challenges related to single color and uneven color distribution. The diffusion models Palette appears severe background coloring while BBDM exhibits a uniform color distribution within the outline but there is a struggle to retain fine details in the generated images. Overall, existing methods encounter challenges in capturing specific intricate details in sketches. In contrast, our proposed method excels in handling complex sketches, leading to images with more realistic visual effects.
Figure 7: Comparison of model training and inference times.
Time Costs. Fig. 7 illustrates the model training and inference times for our model and the nine baselines. Our model exhibits a clear advantage in training time compared to GAN-based methods. Additionally, when compared to the methods based on the Diffusion model, our model demonstrates an inference time improvement of over 100 times. This significant time-saving advantage can greatly benefit designers by reducing costs, improving work efficiency, and increasing productivity.
IV-F Ablation study
We conduct ablation studies to analyze the effectiveness of (i) the Abstract Feature Representation Module (AFRM) and (ii) the Cross-level Skip Connection Module (CSCM).
Quantitative Experiments. As indicated in Table II, the removal of the CSCM between different layers of HAIFIT negatively impacts our model’s performance. CSCM is essential in considering feature correlations across scales and combining different receptive fields, which helps preserve important semantic details. The increase in the PSNR index after removing CSCM is attributed to the loss of sketch detail features, resulting in smoother perceived details. Moreover, removing the AFRM also led to a decline in our model’s performance. AFRM learns long-term dependencies from different perspectives of the sketch using the LSTM module. It takes into account the correlation between adjacent lines in the sketch to increase the details and ensure the consistency of the generated image. Notably, when replacing the LSTM module with the Transformer [28] module, all four performance indicators of the model deteriorated. This decline can be attributed to the temporal dependence offered by the LSTM module, ensuring consistent feature extraction across all directions of the sketch and yielding a more intricate and coherent feature representation. In contrast, the Transformer module, relying on the self-attention mechanism, captures relationships within sequences but lacks strong modeling capabilities in preserving the sequence’s order of features. This outcome underscores the significance of considering the temporal correlation between lines to enhance model performance. Based on the above analysis, the AFRM and CSCM modules enhance our model’s capability to handle complicated sketches. Consequently, HAIFIT enables the generation of high-quality and diverse clothing images while maintaining the design style of the sketches.
TABLE II: The quantitative experimental results on ablation study.
Methods Structural Perceptual
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
w/o AFRM and CSCM 19.7086 0.6324 0.0812 30.5260
w/o AFRM 19.0474 0.6281 0.0821 31.8295
w/o CSCM 19.7189 0.6294 0.0802 29.9711
Transformer 18.7738 0.6259 0.0814 31.1292
HAIFIT (full) 19.0660 0.6338 0.0778 28.5032
Figure 8: Qualitative results of the ablation study on HAIFIT.
Qualitative Experiments. The qualitative results of the ablation study on sketch-to-image generation are presented in Fig. 8. It can be observed that removing the AFRM or CC modules leads to uneven color distribution in the generated images or loss of certain details. Moreover, when both modules are removed simultaneously, these issues become more pronounced. While replacing the Transformer module may also get good results, problems can still be observed in the image color distribution. These findings further confirm the effectiveness of our model.
V Conclusions
In this paper, we introduce HAIFIT, a creative sketch-to-image generation model tailored for fashion design. We effectively enhance both the global consistency of sketch feature vectors and the richness of intricate textures in the latent space through the proposed MFFN and CSCM. This enhancement enables the generation of high-quality, realistic clothing images. Furthermore, We present a fashion clothing dataset HAIFashion for image translation to fill the gap in this field. We conduct extensive experiments comparing our method with state-of-the-art approaches. The qualitative and quantitative results robustly demonstrate the effectiveness of our proposed methodology in preserving sketch details and generating realistic images.